# MobiFuse: Learning Universal Human Mobility Patterns through Cross-domain Data Fusion

### v1

@@ A.0.3 Distribution Refinement
1. Distribution Comparison: Activity frequencies across sub-regions are compared with ground truth data using cosine similarity metrics.
2. Parameter Adjustment: Reference distributions ($D_{md}$, $D_{nmd}$, $\theta$) are iteratively adjusted using: $D^{t+1}=D^{t}+\eta\cdot(F_{target}-F_{current})$ where $F$ represents activity frequency distributions and $\eta$ is a learning rate.
3. Convergence Criteria: The refinement process continues until: $|F_{target}-F_{current}|<\epsilon$ where $\epsilon$ is a predefined threshold.
@@
Appendix B Dataset Details
B.1 Household Travel Surveys
@@
Household Travel Surveys (HTS) have long been a fundamental data source in transportation planning and modeling. These surveys provide comprehensive insights into daily travel patterns by collecting detailed information about individual and household travel behaviors, including complete door-to-door trip chains, activity purposes, timing, and mode choices. This detailed trip chain information makes HTS particularly valuable for understanding human mobility decisions’ sequential nature and underlying patterns.
The training of the mobility pattern reconstruction model requires the use of complete activity chain, including every activity of an agent in a 24-hour day. Two datasets with completed activity chains were utilized in this study to construct the base model.
In order to learn common mobility patterns at the national level, we used the 2017 National Household Travel Survey (NHTS), administered by the Federal Highway Administration (FHWA), to capture nationwide mobility patterns (Administration, 2017). The NHTS includes socio-economic demographics and travel diaries from 129,600 households. We aggregated the original 19 activity types recorded in the NHTS into 15 categories based on the activity locations, which are presented in Table 1. In this study, only the travel diary of each agent in NHTS is used for training.
To learn a specific regional activity pattern, we adopted the activity chain data generated by the Activity-Based Model (ABM) of Southern California Association of Governments (SCAG). The SCAG ABM is a comprehensive modeling system that integrates a series of activity-related choice models, ranging from long-term to short-term decisions, to simulate the activity chains of residents in Southern California. This model generates a synthetic dataset that includes detailed information about the types and sequences of activities, locations, and start/end times. These features make the dataset a valuable tool for analyzing the spatiotemporal dynamics of human activities. Further details of the SCAG ABM and its methodology can be found in  (He et al., 2022; Jiang et al., 2022).
B.2 Time Use Surveys
A key advantage of TUS is its widespread availability across different countries and regions. National statistical offices regularly conduct these surveys as part of their standard data collection efforts, making them more accessible than specialized transportation surveys like HTS. This broad availability is particularly valuable for regions with no HTS. The generalizability of our framework partially comes from the usage of TUS, which can replace the generation’s reliance on HTS with a more accessible data source.
Time Use Surveys (TUS) contain detailed individual-level records of daily activities, where each respondent reports their time allocation alongside their demographic attributes. We leverage this data in two ways: first, the individual records with demographic labels are used to train a demographic classification model, enabling automatic categorization of semantic enriched GPS trajectories into demographic groups. Second, by aggregating these individual records by demographic attributes, we obtain target activity distributions for each demographic group. These aggregated distributions then serve as optimization targets for our demographic adapters optimized by Gradient-Based Distribution (Algorithm 2), allowing the model to adjust activity patterns according to the classified demographic characteristics.
  Output: Optimized adapter collection $\mathcal{A}$
  Process $\mathcal{T}$ to get target distributions $\{\mathbf{D}_{g}\}$
  Initialize $\mathcal{A}=\{\}$
  for $g\in\mathcal{G}$ do
     Initialize adapter weights $\mathbf{a}_{g}$
     repeat
        $\mathbf{Y}=f(\text{input})$
        $\mathbf{P}_{g}^{t}=\text{softmax}(\mathbf{Y}+\mathbf{a}_{g}^{t})$
        $\nabla_{\mathbf{a}_{g}}\mathcal{L}=\mathbf{D}_{g}-\mathbf{P}_{g}^{t}$
        $\mathbf{a}_{g}^{t+1}=\text{clip}(\mathbf{a}_{g}^{t+1},\epsilon,M)$
        if $\|\mathbf{P}_{g}^{t}-\mathbf{D}_{g}\|_{2}\leq\delta$ then
           break
        end if
     until $iteration=max\_iterations$
     $\mathcal{A}=\mathcal{A}\cup\{\mathbf{a}_{g}\}$
  end for
  return $\mathcal{A}$
B.3 GPS and POI Data Integration
@@
These limitations of raw GPS data can be largely addressed through strategic integration with POI data. POI data provides the semantic context necessary to transform anonymous coordinate points into meaningful activity locations. By matching GPS stay points with nearby POIs, we can infer likely activity types and purposes, adding crucial semantic meaning to the raw movement data. Our framework employs Large Language Models (LLMs) for this matching process, enabling more nuanced and context-aware association between locations and activities.
@@
Appendix C Experiment Details
C.1 Los Angeles County Case Study
C.1.1 Data Preparation
Figure 6: Example POI labeling (Liu et al., 2024a)
Figure 7: Example annotated trajectory (Liu et al., 2024a)
To begin our first case studies of LA County, we gathered half-years of GPS points from Veraset, a data vendor. We preprocess the raw GPS trajectories to daily stay points data aggregated by agents, which are consecutive minimum 10-minutes dwell points throughout the GPS trajectory. We then applied filters to acquire around 1 million of relatively high-quality daily stay points data, with a minimum coverage of 55% time of day and more than 1 activity per day per stay points trajectory. Then we fuse the daily stay points and POI data to annotated activity chains using the LLM-based semantic enrichment approach developed in (Liu et al., 2024a). This method combines POI classification using Large Language Models with a Bayesian-based activity inference algorithm to identify the most probable activity type for each stay point. Examples of POI labeling and annotated trajectories are shown in Figure 6 and Figure 7.
@@ C.1.1 Data Preparation
Then we construct a representative synthetic population for Los Angeles County using a rule-based classification approach. Through analysis of our annotated trajectories, approximately 70% contain work-related activities, aligning with LA County’s employment rate (US Bureau of Labor Statistics, 2023). We leverage distinctive activity patterns as classification rules - trajectories containing work activities are classified as belonging to workers, while those with school activities are strongly indicative of young demographics. This rule-based approach, rather than learning-based methods, was deliberately chosen to minimize error propagation in upstream processing stages that could compound into larger biases in downstream tasks. For trajectories where simple rules cannot provide confident classification, we maintain them as unclassified rather than risk misattribution.
C.1.2 Base Model and Its Transferability
@@
The above process prepares semantically enriched but incomplete activity chains for our core reconstruction task. For base model training, we utilize the Southern California Association of Governments (SCAG) Activity-Based Model (ABM) dataset, which provides comprehensive synthetic activity chains for the Los Angeles region. The SCAG ABM integrates various activity-related choice models spanning both long-term and short-term decisions, generating detailed information about activity types, sequences, and temporal patterns (Liao et al., 2024b). We selected this dataset over alternatives like NHTS due to its substantial size - 700,000 samples for training, 200,000 for validation, and 100,000 for testing - which is crucial for effectively training transformer-based architectures.
@@
The training procedure follows a carefully structured three-phase protocol: 1. Initial Warm-up Phase: The model trains on unmasked data for 5 epochs, establishing baseline understanding of complete activity chains 2. Intermediate Phase: Training continues with 40% masked data for 40 epochs, introducing the model to partial data scenarios 3. Final Phase: The remaining epochs use 70% masked activity chains, preparing the model for heavily incomplete real-world data
@@
To ensure robust training, we implement several regularization techniques including dropout, L2 regularization, and early stopping. The model is trained for 120 epochs with a batch size of 512, optimizing a combined loss function that incorporates cross-entropy, transition, and dynamic time warping losses.
The model’s performance in reconstructing activity chains is first evaluated on two survey datasets with ground truth data. Using the test set, 70% of the activity chains are masked and reconstructed. The evaluation measures the similarity between the reconstructed and original activities based on daily activity counts (chain length), activity durations, types, and start/end times.
Algorithm 3 Semi-Supervised Transfer Learning
  Input: Raw activity chain $A_{raw}$, initial model $M^{0}$
  Output: Synthetic activity chain $A_{syn}$
  Generate $A_{0}=M^{0}(A_{raw})$ and initialize $n=1$
  repeat
    Train $M^{n}$ using $A_{n-1}$ with masking
    Generate $A_{n}=M^{n}(A_{raw})$
    Compute JSD similarity $J=\text{JSD}(A_{raw},A_{n})$
    if $J$ has converged then
       $A_{syn}=A_{n}$
       break
    end if
    $n=n+1$
  until convergence or maximum iterations reached
The good performance of the base model $M^{N}$ (trained on NHTS data) is quantified by its values reconstructing masked NHTS data, achieving JSD values between 0.001 and 0.017 in Table 2. However, when applied to the SCAG dataset without transfer learning, the model’s performance declines, with significant JSD values increasing. This indicates that the model trained on national-wide data doesn’t adapt well to regional (Los Angeles) patterns, due to the mobility pattern differences between these two datasets. The JSD values between NHTS and SCAG, ranging from 0.009 to 0.059 across various activity attributes, quantify these differences. For instance, as shown in Figure 4, the activity start time distribution shows a notable peak for SCAG (in red) around 8:00 am, which is less pronounced in the NHTS data (in blue). Moreover, the activity type graph reveals that SCAG data lacks representation in categories like childcare, health care visit, and religious activities, which are present in the NHTS dataset.
To address these discrepancies, we implement transfer learning on model $M^{N}$ (algorithm 3. This approach yields marked improvements, with the SCAG dataset fine-tuned model $M^{S}$ achieving JSD values as low as 0.001 for time and 0.001 for activity type. Figure 4 visually confirms this improvement, showing close alignment between red solid lines and red dash lines across all distributions. These results demonstrate our algorithm’s effectiveness in adapting to different geographical contexts and capturing diverse mobility patterns, despite the initial significant disparities between the SCAG and NHTS datasets.

### v2

@@ 4.3.3 Distribution Refinement
Distribution Comparison: Activity frequencies across sub-regions are compared with ground truth data using cosine similarity metrics.
Parameter Adjustment: Reference distributions ($D_{md}$, $D_{nmd}$, $\theta$) are iteratively adjusted using: $D^{t+1}=D^{t}+\eta\cdot(F_{target}-F_{current})$ where $F$ represents activity frequency distributions and $\eta$ is a learning rate.
Convergence Criteria: The refinement process continues until: $|F_{target}-F_{current}|<\epsilon$ where $\epsilon$ is a predefined threshold.
@@
5 Data Source
5.1 Household Travel Surveys
@@
Household Travel Surveys have long been a fundamental data source in transportation planning and modeling. These surveys provide comprehensive insights into daily travel patterns by collecting detailed information about individual and household travel behaviors, including complete door-to-door trip chains, activity purposes, timing, and mode choices. This detailed trip chain information makes HTS particularly valuable for understanding the sequential nature of human mobility decisions and underlying patterns.
A key contribution of HTS is its standardized categorization of travel activities, which forms the basis for our model’s activity type inference and pattern generation. This standardization provides a consistent vocabulary for describing different types of activities while capturing temporal relationships between activities and their durations. Furthermore, HTS data includes rich sociodemographic information, enabling the understanding of how different population groups engage in various travel behaviors.
In our framework, HTS data serves as the foundation for training our base model, enabling it to learn complex sequential patterns in daily activities and their temporal dependencies. The comprehensive nature of HTS allows our model to capture not only individual activity choices but also the logical connections between successive activities. This learning enables the generation of synthetic activity chains that maintain realistic temporal patterns while respecting logical constraints in human mobility behavior.
However, the limited availability of HTS data poses a significant challenge to mobility modeling, particularly in developing regions. Traditional HTS collection is resource-intensive and costly, making it impractical for many countries to conduct regularly. To address this limitation, our framework implements several innovative approaches. First, through domain adaptation, our model can transfer knowledge from regions with available HTS data to those without, while preserving cultural and regional characteristics. This adaptation process considers regional activity preferences and local transportation contexts.
We further enhance our framework’s capability through a data augmentation approach that generates synthetic travel patterns while maintaining statistical consistency with observed behaviors. This is complemented by our AI network’s novel self-completion mechanism, which can intelligently fill gaps in incomplete travel patterns while maintaining consistency with learned behavioral patterns. The framework also integrates street network information to enhance the realism of generated patterns, ensuring that synthetic travel behaviors reflect realistic constraints of the physical transportation infrastructure.
5.2 GPS and POI Data Integration
@@
These limitations of raw GPS data can be largely addressed through strategic integration with Point of Interest (POI) data. POI data provides the semantic context necessary to transform anonymous coordinate points into meaningful activity locations. By matching GPS stay points with nearby POIs, we can infer likely activity types and purposes, adding crucial semantic meaning to the raw movement data. Our framework employs Large Language Models (LLMs) for this matching process, enabling more nuanced and context-aware association between locations and activities.
@@
5.3 Time Use Survey Data
A key advantage of TUS is its widespread availability across different countries and regions. National statistical offices regularly conduct these surveys as part of their standard data collection efforts, making them more accessible than specialized transportation surveys. This broad availability is particularly valuable for regions where detailed travel surveys are scarce or unavailable.
TUS data is structurally organized by sociodemographic characteristics, providing detailed insights into how different population groups allocate their time. The surveys typically segment data by age, gender, employment status, and other demographic factors, enabling fine-grained analysis of behavioral patterns across different population segments. This demographic sensitivity makes TUS particularly valuable for our framework, as it provides target distributions for activity patterns that vary by population group.
In our framework, these demographic-specific activity distributions serve as the foundation for our demographic adaptation approach. When generating synthetic activity patterns, our model uses TUS-derived distributions as reward signals, encouraging the generation of activities that statistically match observed patterns for each demographic group. This mechanism ensures that generated activities not only follow plausible sequences but also reflect realistic time-use patterns for different population segments.
However, integrating TUS data presents unique challenges. The primary difficulty lies in reconciling the different activity classification systems between TUS and transportation surveys. While HTS focuses on travel-related activities, TUS employs a broader, more general activity classification system. This difference necessitates careful mapping and harmonization of activity categories to maintain consistency in our modeling framework.
The cultural sensitivity of TUS data adds another valuable dimension to our framework. These surveys capture fundamental differences in how societies organize their daily activities - from variations in working hours and meal times to distinct patterns of social interaction and leisure activities. This cultural awareness is crucial for our framework’s ability to adapt to different regional contexts, supporting our goal of generating culturally appropriate mobility patterns across diverse urban environments.
6 Los Angeles (LA) County Case Study
We evaluate our framework through two case studies: Los Angeles (LA) County and Egypt. The LA study demonstrates our complete pipeline capabilities, while Egypt showcases transfer learning in data-scarce regions. Each component in our pipeline - semantic annotation, activity reconstruction, transfer learning (including both activity pattern and demographic adaptation) - is essential in the data fusion process, as removing any component would prevent the generation of complete, demographically representative mobility patterns. We compare the output synthetic data against multiple sources, such as HTS, TUS, and raw GPS points, as our baseline, which helps reflect the disagreement on the same measurements across different data sources.
6.1 Experimental Setup
6.1.1 Data Preprocessing
For our Los Angeles County case study, we gathered half-year GPS points from Veraset (Veraset, 2024), a data vendor. We developed a robust stay point extraction methodology using PySpark-based distributed processing to handle million-scale GPS trajectory analysis efficiently.
Stay Point Detection Algorithm and Validation: We implemented a dual-threshold approach combining temporal (300-second) and spatial (300-meter) criteria with dynamic centroid updating. The algorithm employs haversine distance calculations to maintain spatial accuracy, incrementally updating centroid positions as new points are added to each stay cluster. Sequential processing ensures that genuine activity locations are distinguished from transient GPS noise. Speed-based trajectory filtering using a 30 km/h threshold distinguishes genuine stays from passing trajectories, preventing misclassification of highway travel as activity locations. We implemented H3 hexagonal indexing (resolution 10) for efficient spatial clustering and applied regional boundary validation (33.5-34.5°N, -118.8 to -118.0°W) to ensure LA County geographic focus (Uber Engineering, 2018). Additional quality filters retained only trajectories with minimum 33% time-of-day coverage and more than one activity per day, ensuring adequate daily activity representation.
Our large-scale stay point extraction employs a multi-stage processing pipeline detailed in Algorithm 4. The algorithm processes raw GPS trajectories through temporal clustering, spatial validation, and quality assurance to produce high-fidelity activity locations.
Algorithm 4 Large-Scale Stay Point Extraction Pipeline
1:  Input: Raw GPS trajectories $T$, Temporal threshold $\delta_{t}=300s$, Spatial threshold $\delta_{s}=300m$
2:  Output: Processed stay points $S$ with H3 regional mapping
3:  // Stage 1: Temporal Stay Detection
4:  for each trajectory $t_{i}\in T$ do
5:     $temporal\_stays\leftarrow$ detect_temporal_gaps($t_{i},\delta_{t}$)
6:     $spatial\_stays\leftarrow$ sequential_spatial_clustering($temporal\_stays,\delta_{s}$)
7:  end for
8:  // Stage 2: Regional Filtering and H3 Mapping
9:  $regional\_users\leftarrow$ filter_by_boundary($spatial\_stays$, LA_bounds)
10:  $h3\_stays\leftarrow$ map_to_h3($regional\_users$, resolution=10)
11:  // Stage 3: Speed-based Trajectory Filtering
12:  for each stay sequence $s_{i}\in h3\_stays$ do
13:     if travel_speed($s_{i}$) $>30$ km/h then
14:        classify_as_passing($s_{i}$)
15:     else
16:        classify_as_stay($s_{i}$)
17:     end if
18:  end for
19:  // Stage 4: Quality Assurance
20:  $filtered\_stays\leftarrow$ apply_coverage_filter($h3\_stays$, min_coverage=0.33)
21:  $S\leftarrow$ aggregate_daily_patterns($filtered\_stays$)
22:  return $S$
Following the extraction pipeline, we conducted comprehensive statistical validation of the processed stay points to ensure data quality and behavioral plausibility. The analysis reveals robust spatial and temporal patterns consistent with urban mobility characteristics.
The extracted stay points demonstrate realistic spatial clustering patterns across LA County, with higher density concentrations in urban centers and employment hubs. Geographic density analysis, as shown in Figure 4, exhibits expected urban mobility patterns with concentrated activity in downtown LA, Hollywood, Santa Monica, and major employment corridors, which corresponds with higher population density associated with these areas. In the northern part of LA County, such as Lancaster, the activity density is relatively lower than the population density, because these places are typically underdeveloped residential areas that provide less employment and attractions.
Figure 4: GPS Staypoints and Population Density of LA County
Statistical analysis of stay point characteristics reveals well-distributed activity patterns throughout the week. As shown in Figure 5, the average daily activity frequency shows optimal coverage (mean: 64.0%, median: 61.5%) exceeding our minimum 33% threshold, ensuring adequate temporal representation for reconstruction modeling. Activity duration analysis shows a mean stay duration of 3.6 hours, indicating substantial dwell times consistent with meaningful activity engagement rather than transient GPS noise. Activity type duration distributions follow expected patterns with shorter durations for transitional activities (shopping, meals) and longer durations for anchoring activities (home, work).
Figure 5: Comprehensive Statistical Analysis of processed GPS trajectories
The H3-based spatial clustering effectively captured regional mobility patterns while maintaining computational efficiency. Speed-based filtering successfully distinguished 89% of stay points from passing trajectories, preventing misclassification of highway travel as activity locations. Regional boundary validation ensured 100% geographic consistency within LA County bounds, supporting accurate regional modeling requirements for downstream transportation analysis.
Following preprocessing, we acquired approximately 1 million high-quality daily stay point trajectories, aggregated by individual agents to form complete daily activity patterns suitable for downstream reconstruction modeling.
We fused the daily stay points and POI data following our methodology (Section 4.1.1), which combines POI classification using Large Language Models with a Bayesian-based activity inference algorithm to identify the most probable activity type for each stay point. Examples of POI labeling and annotated trajectories are shown in Figures 6 and 7.
Figure 6: POI labeling using LLM by associating three activity types with likelihood estimation (a) Los Angeles, USA (b) Cairo, Egypt. (Liu et al., 2024)
Figure 7: Semantic annotated trajectories with activity type. (Liu et al., 2024)
@@ 6.1.1 Data Preprocessing
We constructed a representative synthetic population for Los Angeles County using a rule-based classification approach. Through analysis of our annotated trajectories, approximately 70% contain work-related activities, aligning with LA County’s employment rate (US Bureau of Labor Statistics, 2023). We leveraged distinctive activity patterns as classification rules - trajectories containing work activities were classified as belonging to workers, while those with school activities were strongly indicative of young demographics. This rule-based approach, rather than learning-based methods, was deliberately chosen to minimize error propagation in upstream processing stages that could compound into larger biases in downstream tasks. For trajectories where simple rules could not provide confident classification, we maintained them as unclassified rather than risk misattribution.
@@
For base model training, we utilized the Southern California Association of Governments (SCAG) Activity-Based Model (ABM) dataset, which provides comprehensive synthetic activity chains for the Los Angeles region. The SCAG ABM integrates various activity-related choice models spanning both long-term and short-term decisions, generating detailed information about activity types, sequences, and temporal patterns (Liao et al., 2024). We selected this dataset over alternatives like NHTS (National Household Survey Data) (Section 5.1) due to its substantial size - 700,000 samples for training, 200,000 for validation, and 100,000 for testing - which is crucial for effectively training transformer-based architectures.
For demographic adapter training, to establish target distributions $\mathbf{D}_{g}$ for each demographic group as defined in Section 4.2.3, we preprocessed Time Use Survey data at the individual survey level (Section 5.3. For LA County, each TUS entry contains detailed time allocation across daily activities for individual respondents. We mapped activity types from the original TUS classification system to our standardized 15-activity taxonomy used throughout the framework.
Time use data was then aggregated to the demographic group level, computing the percentage of time allocated to each activity category. This preprocessing transforms raw survey responses into probability distributions $\mathbf{D}_{g}\in\mathbb{R}^{|\mathcal{T}|}$ representing typical activity patterns for each of the 12 demographic groups (3 age groups × 2 gender categories × 2 employment statuses). These target distributions serve as targets for the demographic adapter optimization process described in Algorithm 2.
6.1.2 Evaluation Metrics
We employ Jensen-Shannon Divergence (JSD)  (Luca et al., 2021) as our primary metric to evaluate distributional similarities across multiple dimensions: activity chain length, duration, type, and temporal patterns (start/end times), as presented in Equation 11. Lower JSD values indicate better alignment. For spatial validation, we use cosine similarity for Origin-Destination matrices and Mean Absolute Percentage Error (MAPE) for traffic metrics.
$JSD(P\|Q)=\frac{1}{2}\sum_{x\in X}\left[P(x)\log\left(\frac{P(x)}{M(x)}\right)\right]+\frac{1}{2}\sum_{x\in X}\left[Q(x)\log\left(\frac{Q(x)}{M(x)}\right)\right]$ (11)
6.1.3 Training Procedure
The training procedure follows a carefully structured three-phase protocol:
  1. 1.
    Initial Warm-up Phase: The model trains on unmasked data for 10% of total epochs, establishing baseline understanding of complete activity chains.
  2. 2.
    Intermediate Phase: Training continues with 40% masked data for the next 30% of total epochs, introducing the model to partial data scenarios.
  3. 3.
@@ 6.1.3 Training Procedure
    Final Phase: The remaining epochs use 70% masked activity chains, preparing the model for heavily incomplete real-world data.
@@
The training for activity reconstruction (Section 4.1.2), knowledge adapter (Section 4.2.1), and pattern adapter (Section 4.2.2) was conducted on an NVIDIA L40S GPU. As all three tasks share a similar network structure, the batch size was uniformly set to 512, with a learning rate of 0.001. To ensure stable training, we applied regularization techniques including dropout, L2 regularization, and early stopping. The model was trained up to 120 epochs using a combined loss function comprising cross-entropy, transition, and dynamic time warping losses, weighted at 0.7, 0.15, and 0.15, respectively.
The demographic adapter (Section 4.2.3) training was implemented on an NVIDIA RTX 4090 GPU, requiring approximately 10 hours of computation time, averaging 40 minutes per demographic group. Detailed training specifications include: learning rate of 0.01, exponential reward function, and adapter initialization with all weights set to 1.0. Weight constraints maintain a minimum threshold of 0.01 to prevent negative values and apply GradNorm to dynamically handle the loss term weighting, optimizing across 15 activity categories. The training configuration employed a maximum of 1000 epochs for convergence, a batch size of 1500 samples, and a convergence threshold of 0.01 for early stopping. Data processing utilized 16,000 preprocessed activity sequences with 96 time slots (15-minute intervals over 24 hours). The masking strategy applied continuous segment masking with 70% coverage, 8-20 time steps per masked segment (maximum 8 segments), and 60% random activity masking during training. Best model selection was based on the lowest RMSE versus the target distribution from TUS across all training epochs, with automatic early stopping upon convergence detection. This relatively efficient training time, despite the complexity of demographic-specific pattern learning, demonstrates the practical feasibility of our approach for real-world applications.
After training the demographic-aware adapters, we integrated them with the transferred reconstruction model to create a comprehensive system capable of generating both demographically appropriate and temporally coherent activity patterns. This integrated model preserves the base model’s capabilities in temporal pattern recognition and activity dependencies while incorporating demographic sensitivity through the trained adapters.
Using the trained demographic-aware model, we conducted large-scale inference to generate synthetic activity patterns using NVIDIA L40S GPUs. We handled different trajectory data using different masking methods. For trajectories with relatively low loss or high completeness, we could often infer their socio-demographic information. In such cases, the problem became one of completion: we simply matched the corresponding socio-demographic info to the trajectory and used the model to fill in the missing data. However, for data with low completeness, it was not possible to infer any socio-demographic information. We treated this as a generation task, masking as much data as possible—up to 80% of the time in a day—and assigning it a socio-demographic profile. The model then generated a trajectory consistent with this socio-demographic information.
The model generated activity chains through multiple passes - 5 passes for weekdays and 2 passes for weekends - with each pass using different random masks to ensure diversity in the generated patterns. This process allowed the model to explore various possible activity combinations while maintaining demographic consistency. The generation process incorporated both demographic factors and variations in the day of the week to ensure realistic activity patterns between different segments of the population and throughout the week.
6.2 Model Performance Analysis
6.2.1 LLM-based POI Classification and Assignment Validation:
We conducted comprehensive validation using ground truth data from both Los Angeles County and Egypt. Our evaluation reveals the critical advantages of LLM approaches in handling real-world data challenges.
We randomly sampled 500 POI observations from each region and created manual ground truth labels through five independent human annotators following uniform standards. This process enables quantitative assessment of our LLM-based classification accuracy against human judgment.
Using GPT-3.5 as our primary LLM, we achieved 93.4% accuracy and 96.1% F-1 score for Egypt dataset, and 91.4% accuracy with 92.9% F-1 score for LA County dataset. The model’s Hit@1 rates (where the top prediction matches ground truth) reached 74.7% for Egypt and 75.7% for LA, demonstrating consistent performance across diverse geographic and linguistic contexts.
We also validated the POI assignment procedure using our Bayesian-based approach that integrates LLM-derived POI classifications with temporal activity patterns. Using the NHTS California Add-on dataset as ground truth, we created a test dataset of 362 individuals with 2,007 activities, consisting of 1,724 mandatory activities and 283 non-mandatory activities.
Our activity inference validation demonstrates strong overall performance with 91.7% accuracy and 92.3% F-1 score. The model shows particularly robust performance for mandatory activities (home, work, school), achieving 98.3%, 87.2%, and 90.0% accuracy, respectively, validating the LLM’s ability to classify POIs associated with these fundamental daily activities correctly. Non-mandatory activities show more varied performance, with categories like healthcare (100% accuracy) and religious activities (100% accuracy) performing excellently, while more ambiguous activities like "visit friends" (52.2% accuracy) prove more challenging due to their context-dependent nature.
To assess real-world robustness, we conducted comprehensive noise tolerance testing by introducing Gaussian noise at three levels (5m, 10m, and 20m standard deviation) to simulate GPS inaccuracies common in urban environments. The model maintains strong stability across noise levels, with Accuracy@3 remaining above 84% even at 20m noise levels (88.4% at 5m, 85.3% at 10m, 84.2% at 20m), demonstrating the framework’s resilience to typical GPS positioning errors encountered in real-world deployment scenarios.
The Bayesian inference component effectively combines LLM-derived POI activity probabilities with temporal likelihood distributions derived from survey data, enabling context-aware activity classification. This dual validation approach—both direct POI classification and end-to-end activity inference—provides comprehensive evidence of the LLM framework’s effectiveness in handling the complete semantic trajectory annotation pipeline from raw GPS coordinates to meaningful activity sequences.
For baseline comparison, we implemented a rule-based heuristic POI classification method that maps OSM tags such as amenity, shop, tourism, leisure, etc., to the 15 activity type codes using a fixed priority and mapping scheme. This approach assigns each POI up to three most likely activity types based on the presence and priority of OSM tags, and evaluates performance using top-1, top-2, and top-3 accuracy as well as the weighted F1 score. As shown in Table 2, the heuristic method achieves an Accuracy@1 of 64.3% for Egypt and 49.4% for LA County, which is substantially lower than the LLM-based approaches (gpt-3.5 and gpt-4). The heuristic’s Accuracy@2 and @3 are also much lower, indicating limited ability to capture ambiguous or multi-purpose POIs. The F1 scores further highlight the gap, with the heuristic baseline reaching only 65.7% (Egypt) and 51.7% (LA), compared to over 90% for LLMs. These results demonstrate that while the heuristic method provides a transparent and interpretable baseline, it is significantly outperformed by LLM-based classification, especially in handling complex or ambiguous POI categories.
Table 2: Performance metrics for POI classification with different approaches.
Metric Egypt (gpt3.5) Egypt (gpt4) Egypt (Heuristic) LA (gpt3.5) LA (gpt4) LA (Heuristic)
Accuracy@1 90.3% 93.4% 64.3% 82.5% 91.4% 49.4%
Accuracy@2 22.6% 9.7% 4.4% 15.4% 8.7% 0.4%
Accuracy@3 6.1% 9.0% 0.0% 1.9% 7.0% 0.0%
F-1 Score 91.5% 96.1% 65.7% 82.1% 92.9% 51.7%
6.2.2 Base Model Evaluation on Survey Data
The performance of the base model $M^{N}$ in reconstructing 70% masked NHTS data and its transferability to the SCAG datasets is quantified in Table 3. The model achieves strong performance on NHTS data, with JSD values ranging from 0.001 to 0.017 across all metrics. However, directly applying the model to SCAG data results in performance degradation indicated by increasing JSD value, revealing significant regional differences in mobility patterns. The JSD values between NHTS and SCAG, which range from 0.009 to 0.059 across various activity attributes, quantify these differences and are visualized in Figure 8. Supervised transfer learning is applied to $M^{N}$ to obtain $M^{S}$, yielding low JSD values when reconstructing the SCAG dataset. This result shows the algorithm effectively adapts to different geographical contexts and data formats, despite initial disparities between SCAG and NHTS datasets.